# Open-Set Gait Recognition from Sparse mmWave Radar Point Clouds

### v2

@@ VI-C Ablation studies
V1: In this version, the class centroids $\boldsymbol{\mu}_{i}$ defined in Section IV-B are learnt by the network instead of being fixed. More specifically, we replace the deterministic mapping $f$ with a multi-layer perceptron with $3$ hidden layers of dimension $(16,32,64)$.
@@ VII Concluding Remarks
To address the lack of publicly available high-quality data for this task, we collect and publicly release mmGait10, a dataset containing radar point cloud traces from $10$ human subjects, each with three distinct walking modalities. Using mmGait10, we compare our approach to the closest existing state-of-the-art network for OSGR, OR-CED. Our results show the superiority of our approach in terms of OSGR F1111 score in different settings and openness levels. We additionally validate each architectural choice, by performing an extensive ablation study involving three variants of PCAA. Our ablation studies clearly show that the architectural features of PCAA, tailored for point cloud input sequences, are crucial in our setting. This highlights the necessity of specifically designing architecture blocks and learning strategies to deal with complex data structures like radar point clouds.

### v4

@@ V-C Ablation studies
V1: In this version, the class centroids $\boldsymbol{\mu}_{i}$ defined in Section III-B are learnt by the network instead of being fixed. More specifically, we replace the deterministic mapping $f$ with a multi-layer perceptron with $3$ hidden layers of dimension $(16,32,64)$.
@@ VI Concluding Remarks
To address the lack of publicly available high-quality data for this task, we collect and publicly release mmGait10, a radar dataset containing both point cloud traces and $\mu$D spectrograms of $10$ human subjects, each walking with three distinct walking modalities. Using mmGait10, we compare our approach to the most recent existing state-of-the-art approaches for OSGR, OR-CED and L-GM. To provide a comprehensive and fair comparison, we evaluated the performance of the baselines both on $\mu$D spectrograms (the data type they were designed for) and on point cloud data, which is the focus of this work.
Our results show the superiority of our approach in terms of OSGR F1111 score in different settings and openness levels, clearly confirming that a simple adaptation of existing gait recognition strategies for $\mu$D signatures is not sufficient to achieve the same performance when using sparse and noisy point clouds. We additionally validate each architectural choice, by performing an extensive ablation study involving three variants of PCAA. Our ablation studies clearly show that the architectural features of PCAA, tailored for point cloud input sequences, are crucial in our setting. This highlights the necessity of specifically designing architecture blocks and learning strategies to deal with complex data structures like spatio-temporal radar point clouds.